OpenStudy (anonymous):

Solve equation: 1/2lg(3x+1)-1=lg(x-3)-lg(5)

OpenStudy (anonymous):

What is lg ?

OpenStudy (anonymous):

natural log

OpenStudy (anonymous):

Just wright it as ln.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

ok, ok

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\ln(3x+1)-1=\ln(x-3)-\ln(5)\] \[\ln(\sqrt{3x+1})-\ln(x-3)=1-\ln(5)\] \[\ln(\frac{ \sqrt{3x+1} }{x-3})=\ln(e)-\ln(5)\] \[\ln(\frac{ \sqrt{3x+1} }{x-3})=\ln(e)-\ln(5)\] \[\ln(\frac{ \sqrt{3x+1} }{x-3})=\ln(e)-\ln(5)\]I am losing my connection will finish this